# Reconstructing Galaxy Cluster Mass Maps using Score-based Generative Modeling

### v1

@@ Abstract
We present a novel approach to reconstruct gas and dark matter projected density maps of galaxy clusters using score-based generative modeling. Our diffusion model takes in mock SZ and X-ray images as conditional inputs, and generates realizations of corresponding gas and dark matter maps by sampling from a learned data posterior. We train and validate the performance of our model by using mock data from a hydrodynamical cosmological simulation. The model accurately reconstructs both the mean and spread of the radial density profiles in the spatial domain to within 5%, indicating that the model is able to distinguish between clusters of different sizes. In the spectral domain, the model achieves close-to-unity values for the bias and cross-correlation coefficients, indicating that the model can accurately probe cluster structures on both large and small scales. Our experiments demonstrate the ability of score models to learn a strong, nonlinear, and unbiased mapping between input observables and fundamental density distributions of galaxy clusters. These diffusion models can be further fine-tuned and generalized to not only take in additional observables as inputs, but also real observations and predict unknown density distributions of galaxy clusters.
journal: The Open Journal of Astrophysics
@@ 1 Introduction
Galaxy clusters are the largest gravitationally bound systems in the universe, typically comprised of hundreds of galaxies and a hot intracluster medium (ICM) embedded in a dark matter halo. The cluster mass scales with galaxy velocity dispersion, a relation famously used to postulate the presence of dark matter in the Coma cluster (Zwicky, 1933). The number density of galaxy clusters as a function of their mass is a cosmological probe, and is sensitive to the distribution and evolution of large-scale structures in the cosmic web. See Allen et al. (2011) and Kravtsov & Borgani (2012) for reviews.
@@
Concurrently, image-to-image deep learning models have been used to effectively model galaxy cluster structures by learning a compressed latent space representation and sampling from a learned distribution. These generative models compress the input into a latent space by extracting the most useful information that is further used to sample outputs, capturing and appropriately modeling the intrinsic uncertainty in incomplete observations. Examples of reconstruction from the latent space include random sampling of hyper-realistic large cosmological structures (e.g. Ullmo et al., 2021), and accurate reconstruction of galaxy images (e.g. Schawinski et al., 2018) and large-scale structures (e.g. Rodriguez et al., 2018; Perraudin et al., 2019). Other interesting applications of image generation include reconstructing dark matter maps from cluster observables (e.g. de Andres et al., 2024), classification of galaxy mergers (e.g. Arendt et al., 2024), denoising and deconvolving galaxy images (e.g. Schawinski et al., 2017; Hemmati et al., 2022) and increasing the resolution of cosmological simulations (e.g. Li et al., 2021).
@@
A recent paradigm shift in image-to-image generative modeling is the development of diffusion or score-based generative modeling, which is a probabilistic framework that reformulates the previous sampling processes from learned data distributions. By increasingly injecting noise into our data and learning the inverse denoising procedure, we can map a prior distribution to the data distribution and subsequently sample from it. Score matching with Langevin Dynamics (SMLD, e.g. Vincent, 2011; Song & Ermon, 2019; Song et al., 2021) learns the score, or the gradient of the log probability density, of the data distribution which is used by Langevin dynamics to generate data using stochastic differential equations (SDEs). An analogous score-based framework is Denoising Diffusion Probabilistic Modeling (DDPM, e.g. Sohl-Dickstein et al., 2015; Ho et al., 2020) which implicitly learns the score by training a sequence of models to reverse the noise injection. In the context of astrophysics generative modeling, this probabilistic framework can learn and sample images from posterior distributions rather than points estimates: recent work in astrophysics has been applied to a variety of tasks, including sampling galaxy spectra (e.g. Doorenbos et al., 2022, 2024), super-resolution gravitational lensing maps (e.g. Reddy et al., 2024), satellite galaxies and subhalo populations (e.g. Nguyen et al., 2024; Bourdin et al., 2024), and initial conditions of the universe (e.g. Legin et al., 2023).
@@
We present a score-based generative model that learns the predictive posterior distribution of gas and dark matter maps of galaxy clusters. This powerful probabilistic modeling allows us to subsequently sample realizations of these maps conditioned on SZ and X-ray inputs. In Section 2, we detail the process of generating mock observations from a large hydrodynamical simulation as training and testing data for our models. In section 3, we describe the diffusion model pipeline and elaborate on the statistical framework of learning the score of the data distribution. Finally, in section 4 we present the estimated gas and dark matter maps, the reconstructed accuracy of the density profiles in the spatial domain, and the bias and cross correlation coefficients in the spectral domain.
@@ 2.1 Cosmological Simulation
Many supervised machine learning models, such as the image-to-image diffusion model that will be described in Section 3, require large amounts of data for training. We generate a large number of simulated galaxy clusters by running a cosmological simulation of galaxy clusters with an updated version of HYPER (He et al., 2022), a hydro-particle code for efficient and rapid simulations of baryons and dark matter. The hydro simulation used here has a moderately large volume so as to produce many galaxy clusters, paired with a high enough resolution to model internal cluster structures. This combination of large volume and high resolution is made possible due to the novel approach of subgrid models for the thermodynamics of the ICM and IGM, which results in three orders of magnitude speedup compared to a standard hydrodynamics code. In particular, we use an ICM pressure profile and IGM temperature-density relation, both of which can be systematically varied for parameter space studies, and these facilitate the creation of mock X-ray and SZ observations from the simulated ICM.
@@
For the gravity solver, the particle-mesh (PM) algorithm is upgraded to particle-particle-mesh (P3M; Trac et al., 2015) for improved spatial resolution. For the hydro solver, some components of the hydro-particle-mesh (HPM; Gnedin & Hui, 1998; He et al., 2022) algorithm are replaced with elements from smoothed particle hydrodynamics (SPH; see Springel, 2010, for a review). Particle densities and pressure gradients are not calculated using a mesh, but rather by summing over and weighting particles within the individual smoothing lengths. The kick-drift-kick leapfrog integrator now uses adaptive time steps for improved efficiency. The subgrid model for the intracluster medium (ICM) is based on the debiased pressure profile (DPP; He et al., 2021), which adjusts for hydrostatic mass bias by combining results from X-ray observations (e.g. Böhringer et al., 2007) with cosmological simulations (e.g. Barnes et al., 2021).
@@
The simulation is based on the following cosmological parameters: $\Omega_{\mathrm{b}}=0.045$, $\Omega_{\mathrm{m}}=0.3$, $\Omega_{\Lambda}=0.7$, $h=0.7$, $\sigma_{8}=0.8$, and $n_{\mathrm{s}}=0.96$. There are $N_{\mathrm{dm}}=1024^{3}$ dark matter particles and $N_{\mathrm{gas}}=1024^{3}$ gas particles in a comoving box of side length $L_{\mathrm{sim}}=500\ h^{-1}\mathrm{Mpc}$. The particle masses are $m_{\mathrm{dm}}=1.18\times 10^{10}\ M_{\odot}$ and $m_{\mathrm{gas}}=2.08\times 10^{9}\ M_{\odot}$, and the gravitational softening length is $\epsilon=44\ \mathrm{kpc}$. A spherical overdensity halo finder is used to locate halos with density equal to 200 times the average matter density. The halo mass function is accurate to $\lesssim 5\%$ for halos with $\gtrsim 200$ (gas and dark matter) particles. The simulation was run on the NASA Endeavour supercomputer using 256 cores and taking only 9000 CPU-hours.
@@ 2.2 Galaxy Cluster Sample
The HYPER simulation described above produces 6,233 galaxy clusters at redshift $z=0$. The smallest halo has mass $M_{200a}=1.00\times 10^{14}\ M_{\odot}$ and is resolved with $\sim 14000$ particles within a radius $R_{200a}=1.44\ \mathrm{Mpc}$. The largest halo has $M_{200a}=4.82\times 10^{15}\ M_{\odot}$ and contains $\sim 700,000$ particles within $R_{200a}=5.21\ \mathrm{Mpc}$. Figure 1 shows that the number of halos decreases exponentially in the high-mass tail of the distribution. To avoid biases in the ML model, we create a flat distribution with logarithmic mass (e.g. Ntampaka et al., 2015; Ho et al., 2019). For the higher-mass clusters, we generate more sightlines by rotating the particle coordinates. For the lower-mass clusters, we subsample if necessary.
For a given cluster sightline, we first construct three-dimensional fields such as density and pressure by mapping the particles onto a Cartesian mesh with side length of 12.8 Mpc, which is slightly larger than the diameter of the largest halo. For a given field, we then project along up to three cardinal axes to make two-dimensional images with pixel size $l_{\mathrm{pixel}}=100\ \mathrm{kpc}$ and $N_{\mathrm{pixel}}=128$ pixels per side length. The pixel size is approximately twice the gravitational softening length and comparable to the effective resolution of the HYPER simulation. The image resolution is comparable to current X-ray observations and to followup SZ observations. Figure 1 shows that the number of images generated is flat with approximately 3000 per 0.1 dex in mass, except at the very highest masses where we impose a maximum number of sightlines per cluster.
Following the SPH approach, the density field for particle species $\mathrm{x}=\mathrm{Gas,DM}$ is calculated using a weighted summation over nearby particles $j$ with mass $m_{j}$ and smoothing length $h_{j}$:
@@
where $q=r/h\subset[0,1]$, consistent with that used in the HYPER simulation. Other fields like the gas pressure are calculated similarly to Eq. 1 (see Springel, 2010, for a review).
We calculate the normalized projected density for component $\mathrm{x}=\mathrm{Gas,DM}$ by integrating the density field along a given direction:
@@
where $\bar{\rho}_{\mathrm{x}}=\Omega_{\mathrm{x}}\rho_{\mathrm{crit}}$ is the cosmic average density. We make up to three projected density images for a given density field by summing the array $\rho_{\mathrm{x}}[i,j,k]$ along each of the three cardinal axes and multiplying by the pixel length.
@@ 2.3 Sunyaev-Zel’dovich Effect
The SZ effect is commonly considered to have two main components. The thermal SZ (TSZ) effect arises from inverse Compton scattering of the CMB with hot electrons, while the kinetic SZ (KSZ) effect is a Doppler term due to scattering with fast electrons. In this paper, we are interested in the TSZ effect from the ICM. The CMB temperature fluctuations $\Delta T/T_{\mathrm{cmb}}$ are related to the Compton $y$ parameter multiplied by a frequency-dependent function. The Compton $y$ parameter is proportional to the integrated electron pressure along a given sightline,
@@
For a given electron pressure field $P_{\mathrm{e}}(\bm{r})$, we make up to three Compton $y$ images with signal,
@@
by summing the pressure array along each of the three cardinal axes and multiplying by the pixel length. While the Compton $y$ parameter is dimensionless, it is often multiplied by $T_{\mathrm{cmb}}$ to have convenient units of $\mu$K.
@@ 2.4 X-ray Emission
where $g_{\mathrm{ff}}$ is the free-free gaunt factor, $T$ is the plasma temperature, $n_{\mathrm{e}}$ is the electron number density, and $z_{i}$ and $n_{i}$ are the charge and number density of the various ions $i$ (see Sarazin, 1988, for a review). For a given luminosity density field $l_{\mathrm{bol}}(\bm{r})$, we make up to three X-ray surface brightness images with signal
@@ 3 Machine Learning
Our deep learning pipeline consists of two stages. First, we pre-process the mock observations and split our datasets into training and testing portions (Section 3.1). Second, we train the score model on the training dataset to reconstruct the gas and dark matter density maps conditioned on the SZ and X-ray inputs. In Section 3.2, we describe the framework of stochatic differential equations and how score models utilize Langevin dynamics sampling to learn the data distribution.
@@ 3.1 Dataset Construction
Figure 2 shows sample simulation maps of gas and dark matter, and mock observations of SZ and X-ray maps of a cluster. All images are 128 by 128 pixels, with 0.1 Mpc/pixel or 12.8 Mpc in physical length. Each type of image is normalized to within 0 and 1 with the same bounds, so as to not cause large gradient updates within the network weights. Note that this operation is global: we want to preserve the relative intensities between clusters of different sizes, since we want our model to generate the mass maps with intensities associated with the input galaxy cluster size.

### v2

@@ Abstract
We present a novel approach to reconstruct gas and dark matter projected density maps of galaxy clusters using score-based generative modeling. Our diffusion model takes in mock SZ and X-ray images as conditional inputs, and generates realizations of corresponding gas and dark matter maps by sampling from a learned data posterior. We train and validate the performance of our model by using mock data from a cosmological simulation. The model accurately reconstructs both the mean and spread of the radial density profiles in the spatial domain, indicating that the model is able to distinguish between clusters of different mass sizes. In the spectral domain, the model achieves close-to-unity values for the bias and cross-correlation coefficients, indicating that the model can accurately probe cluster structures on both large and small scales. Our experiments demonstrate the ability of score models to learn a strong, nonlinear, and unbiased mapping between input observables and fundamental density distributions of galaxy clusters. These diffusion models can be further fine-tuned and generalized to not only take in additional observables as inputs, but also real observations and predict unknown density distributions of galaxy clusters.
@@ 1 Introduction
Galaxy clusters are the largest gravitationally bound systems in the universe, typically comprised of hundreds of galaxies and a hot intracluster medium (ICM) embedded in a dark matter halo. The cluster mass scales with galaxy velocity dispersion, a relation famously used to postulate the presence of dark matter in the Coma cluster (Zwicky, 1933). The number density of galaxy clusters as a function of their mass is a cosmological probe, and is sensitive to the distribution and evolution of large-scale structures, as density is a probe into the growth of structure which depends on cosmological parameters (see Allen et al. (2011) and Kravtsov & Borgani (2012) for reviews).
@@
Concurrently, image-to-image deep learning models, notably autoencoder-style networks such as Generative Adversarial Models (GANs) and Variational Autoencoders (VAEs), are very good at solving ill-posed astrophysics problems that require learning a compressed latent space representation and sampling from a learned distribution: these models learn to capture and appropriately model the intrinsic uncertainty in incomplete observations. Examples of reconstruction from the latent space include random sampling of hyper-realistic large cosmological structures (e.g. Ullmo et al., 2021), and accurate reconstruction of galaxy images (e.g. Schawinski et al., 2018) and large-scale structures (e.g. Rodriguez et al., 2018; Perraudin et al., 2019). Other interesting applications of image generation include reconstructing total projected matter maps from cluster observables (e.g. de Andres et al., 2024), classification of galaxy mergers (e.g. Arendt et al., 2024), denoising and deconvolving galaxy images (e.g. Schawinski et al., 2017; Hemmati et al., 2022) and increasing the resolution of cosmological simulations (e.g. Li et al., 2021).
@@
A recent paradigm shift in image-to-image generative modeling is the development of diffusion or score-based generative models (SGMs), which is a probabilistic framework that reformulates the previous sampling processes from learned data distributions. By increasingly injecting noise into our data and learning the inverse denoising procedure, we can map a prior distribution to the data distribution and subsequently sample from it. Score matching with Langevin Dynamics (SMLD, e.g. Vincent, 2011; Song & Ermon, 2019; Song et al., 2021) learns the score, or the gradient of the log probability density, of the data distribution which is used by Langevin dynamics to generate data using stochastic differential equations (SDEs). An analogous score-based framework is Denoising Diffusion Probabilistic Modeling (DDPM, e.g. Sohl-Dickstein et al., 2015; Ho et al., 2020) which implicitly learns the score by training a sequence of models to reverse the noise injection. SGMs offer many advantages over traditional GANs and VAEs: not only are SGMs less prone to mode collapse to an average solution during training, they also learn the noise of the problem at all scales in a structured manner, which is crucial for astronomical data that may be noisy, incomplete, or lossy. SGMs also optimize using an explicit likelihood surrogate, whose stability scales well with dimension, as opposed to the amortized inference procedure of VAEs that can under-represent the posterior, or the adversarial losses used by GANs that is known to be unstable. In the context of astrophysics generative modeling, diffusion models have been able to robustly sample various posterior distributions: recent work in astrophysics has been applied to a variety of tasks, including sampling galaxy spectra (e.g. Doorenbos et al., 2022, 2024), super-resolution gravitational lensing maps (e.g. Reddy et al., 2024), satellite galaxies and subhalo populations (e.g. Nguyen et al., 2024; Bourdin et al., 2024), and initial conditions of the universe (e.g. Legin et al., 2023).
@@
We present a score-based generative model that learns the predictive posterior distribution of gas and dark matter maps of galaxy clusters. This powerful probabilistic model allows us to subsequently sample realizations of these maps conditioned on SZ and X-ray inputs. In Section 2, we detail the process of generating mock observations from a cosmological simulation as training and testing data for our models. In section 3, we describe the diffusion model pipeline and elaborate on the statistical framework of learning the score of the data distribution. Finally, in section 4 we present the estimated gas and dark matter maps, the reconstructed accuracy of the density profiles in the spatial domain, and the bias and cross correlation coefficients in the spectral domain.
@@ 2.1 Cosmological Simulation
Many supervised machine learning models, such as the image-to-image diffusion model that will be described in Section 3, require large amounts of data for training. We generate a large number of simulated galaxy clusters by running a cosmological simulation with an updated version of HYPER (He et al., 2022), a hydro-particle code for efficient and rapid simulations of baryons and dark matter. The hydro simulation used here has a moderately large volume so as to produce many galaxy clusters, paired with a high enough resolution to model internal cluster structures. This combination of large volume and high resolution is made possible due to the novel use of subgrid models for the thermodynamics of the ICM and intergalactic medium (IGM), resulting in a three-order-of-magnitude speedup compared to a standard hydrodynamics code.
@@
We summarize the recent updates to HYPER, which will be presented in more detail in an upcoming paper. For the gravity solver, the particle-mesh (PM) algorithm is upgraded to particle-particle-mesh (P3M; Trac et al., 2015) for improved spatial resolution. For the hydro solver, some components of the hydro-particle-mesh (HPM; Gnedin & Hui, 1998; He et al., 2022) algorithm are replaced with elements from smoothed particle hydrodynamics (SPH; see Springel, 2010, for a review). Particle densities and pressure gradients are not calculated using a mesh, but rather by summing over and weighting particles within the individual smoothing lengths. The kick-drift-kick leapfrog integrator now uses adaptive time steps for improved efficiency.
To generate mock SZ and X-ray observations of galaxy clusters, we simulate the ICM using subgrid models. These models include a dark matter halo model, an ICM pressure profile, and an IGM temperature-density relation, all of which can be systematically varied for parameter space studies. In particular, the ICM pressure profile is based on the debiased pressure profile (DPP; He et al., 2021), which accounts for hydrostatic mass bias by combining results from X-ray observations (e.g. Böhringer et al., 2007) with cosmological simulations (e.g. Barnes et al., 2021). In He et al. (2022), the HYPER simulation results agree well with the halo model expectations for the density, temperature, and pressure radial profiles. Additional details are provided in Appendix A.
@@
The simulation is based on the following cosmological parameters: $\Omega_{\mathrm{b}}=0.045$, $\Omega_{\mathrm{m}}=0.3$, $\Omega_{\Lambda}=0.7$, $h=0.7$, $\sigma_{8}=0.8$, and $n_{\mathrm{s}}=0.96$. There are $N_{\mathrm{dm}}=1024^{3}$ dark matter particles and $N_{\mathrm{gas}}=1024^{3}$ gas particles in a comoving box of side length $L_{\mathrm{sim}}=500\ h^{-1}\mathrm{Mpc}$. The particle masses are $m_{\mathrm{dm}}=1.18\times 10^{10}\ M_{\odot}$ and $m_{\mathrm{gas}}=2.08\times 10^{9}\ M_{\odot}$, and the gravitational softening length is $\epsilon=44\ \mathrm{kpc}$. A spherical overdensity halo finder is used to locate halos, where the mass $M_{\mathrm{200m}}$ within radius $R_{\mathrm{200m}}$ has an average density that is equal to 200 times the matter density. The halo mass function is accurate to $\lesssim 5\%$ for halos with $\gtrsim 200$ (gas and dark matter) particles. The simulation was run on the NASA Endeavour supercomputer using 256 cores and taking only 10k CPU-hours.
@@ 2.2 Galaxy Cluster Sample
The HYPER simulation described above produces 6,233 galaxy clusters at redshift $z=0$. The smallest halo has mass $M_{\mathrm{200m}}=1.00\times 10^{14}\ M_{\odot}$ and is resolved with $\sim 14000$ particles within a radius $R_{\mathrm{200m}}=1.44\ \mathrm{Mpc}$. The largest halo has $M_{\mathrm{200m}}=4.82\times 10^{15}\ M_{\odot}$ and contains $\sim 700,000$ particles within $R_{\mathrm{200m}}=5.21\ \mathrm{Mpc}$. Figure 1 shows that the number of halos decreases exponentially in the high-mass tail of the distribution. To avoid biases in the ML model, we create a flat distribution with logarithmic mass (e.g. Ntampaka et al., 2015; Ho et al., 2019). Overabundant low-mass halos are subsampled, while additional sightlines are generated for underabundant high-mass halos. Sightlines are sampled from a Fibonacci sphere (e.g. González, 2010) to ensure a more uniform and isotropic distribution. To generate each sightline, the particle coordinates are rotated so that the original $z$-axis aligns with the specified direction. Figure 1 shows that the number of cluster images generated is flat with approximately 3000 per 0.1 dex in mass, except at the very highest masses where we impose a maximum number of sightlines per cluster.
For a given cluster sightline, we first construct three-dimensional fields, such as density and pressure, by mapping the particles onto a Cartesian mesh. The mesh length $l_{\mathrm{mesh}}=12.8$ Mpc is slightly larger than the diameter of the largest halo, while the cell size $l_{\mathrm{cell}}=100$ kpc is approximately twice the gravitational softening length and comparable to the effective resolution of the HYPER simulation. Following the SPH approach, the density field for each particle species $\mathrm{x}=\mathrm{Gas,DM}$ is calculated using a weighted summation over nearby particles $j$ with mass $m_{j}$ and smoothing length $h_{j}$:
@@
where $q=r/h\in[0,1]$, consistent with the implementation used to compute densities for both gas and dark matter particles in the HYPER simulation. Each particle has an individual smoothing length that is adaptively calculated during the simulation. Note that this smoothing procedure differs from the gravitational softening, which uses a spline kernel with a constant softening length. Other fields such as the electron pressure (Section 2.3) and X-ray luminosity density (Section 2.4) are calculated similarly to Eq. 1 (see Springel, 2010, for a review).
For a given field, we then project along each of the three cardinal axes to make two-dimensional images with $N_{\mathrm{pixel}}=128$ pixels per side. The image resolution, set by the pixel size $l_{\mathrm{pixel}}=100\ \mathrm{kpc}$, is comparable to current X-ray observations and to followup SZ observations. For a given density field $\rho_{\mathrm{x}}(\bm{r})$, we make up to three projected density images with signal,
@@
by summing the density array $\rho_{\mathrm{x}}[i,j,k]$ along each of the three cardinal axes and multiplying by the pixel length. The projected density images are normalized using the cosmic average density $\bar{\rho}_{\mathrm{x}}=\Omega_{\mathrm{x}}\rho_{\mathrm{crit}}$, making them dimensionless.
@@ 2.3 Sunyaev-Zel’dovich Effect
We focus on the thermal SZ (tSZ) effect, which arises from inverse Compton scattering of the CMB with hot electrons. The CMB temperature fluctuations $\Delta T/T_{\mathrm{cmb}}$ are related to the Compton $y$ parameter multiplied by a frequency-dependent function. The Compton $y$ parameter is proportional to the integrated electron pressure along a given sightline,
@@
In the HYPER simulation at $z=0$, both the ICM and IGM are modeled as fully ionized, with one electron per hydrogen and two electrons per helium. A mean molecular weight $\mu_{\mathrm{e}}=1.13$ is used when calculating the electron density or pressure. For a given electron pressure field $P_{\mathrm{e}}(\bm{r})$, we make up to three Compton $y$ images with signal,
@@
by summing the pressure array along each of the three cardinal axes. While the Compton $y$ parameter is dimensionless, it is often multiplied by $T_{\mathrm{cmb}}$ to have convenient units of $\mu$K.
@@ 2.4 X-ray Emission
where $g_{\mathrm{ff}}$ is the free-free gaunt factor, $T$ is the plasma temperature, $n_{\mathrm{e}}$ is the electron number density, and $z_{i}$ and $n_{i}$ are the charge and number density of the various ions $i$ (see Sarazin, 1988, for a review). In the HYPER simulation, the ICM is modeled as a fully ionized plasma with an average metallicity of $0.3Z_{\odot}$ at $z=0$. For a given luminosity density field $l_{\mathrm{bol}}(\bm{r})$, we make up to three X-ray surface brightness images with signal,
@@
For this pilot study, we omit noise from the mock SZ and X-ray images to assess the baseline performance of the method. Realistic noise models will be incorporated in future work before applying the approach to observational data.
@@ 3 Machine Learning
Our deep learning pipeline consists of two stages. First, we pre-process the mock observations and split our datasets into training and testing portions (Section 3.1). Second, we train the score model on the training dataset to reconstruct the gas and dark matter density maps conditioned on the SZ and X-ray inputs. In Section 3.2, we describe the framework of stochastic differential equations and how score models utilize Langevin dynamics sampling to estimate the data distribution.
@@ 3.1 Dataset Construction
Figure 2 shows sample simulation maps of gas and dark matter, and mock observations of SZ and X-ray maps of a cluster. All images are 128 by 128 pixels, with 0.1 Mpc/pixel or 12.8 Mpc in physical length. We normalize these images to within 0 and 1 using global upper and lower bounds for each type of image since we want to preserve the relative intensities between clusters of different sizes: we want our model to generate the mass maps with intensities associated with the input galaxy cluster size.